PREX/CREX Oct 31, 2019 14:00

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  373 507 581
 Bluejeans link: https://bluejeans.com/373507581
 Offline room:   CEBAF Center F224-225

Agenda

  1. Ongoing CREX prep activities: all
    • DM: working to work on the wiki pages for the run plans
    1. Schedule: Don
      • Don: improptu meeting regarding the installation and beam in the hall and schedule. The end of next week we will install the Ca target so we can't interrupt the vacuum after that. The SAM installation needs to happen by the end of next Th. Techs will be ready to install next Wed morning 9AM. First beam in the hall Nov 13th.
      • KP: the alignment has been done. CG: can someone track the results?
      • DM: why is the target going in so early? BM: not clear. SC: target cooldown is scheduled for Nov 11th.
    2. SAM cans: Ciprian CREX-Install-Album
      • CG: manufacturing complete; welding to be done today and tomorrow; shipping Monday afternoon.
      • KP: we should do some basic QA. CG: will do;
  2. BCM testing: Cameron, supplemental evidence: see page 100 of slug 70 PQB plots
    • CC: will work further on the BCM receivers.
    • slide 11: CP: the anticorrelation could indicate that the average of the analogs would be useful for normalization. KP: this is part of the intermittend noise which appeared some times.
    • KP: for the analog noise problem did you show that it's repeatble when going on an off line-sync? CC: yes; will try to repeat the test. KP: the suggestion is that we were sensitive to 60Hz (or multiple) during the PREX2 run.
  3. PREX2 analysis status: Paul K
    • PK: working with the WACs to recheck the run lists and we are almost ready to run the respin.
    • PK: we'll start refocusing on development for CREX soon.
  4. Other
    1. SC: tested for several days with water flowing and we didn't see any issues. Vacuum sat at 1e-6.
    2. SC: didn't take any further pictures of the ladder. The Ca target had a raster mark that was centered.
    3. CC: should we have a meeting for the transverse data? KP: it would be useful to have a discussion but that is in Ryan's and Robert's hands. We should focus on the main analysis for now. PK: let's revist this after the respin is started.

Present

In F245

Bob M, Sanghwa P, Ciprian G, Amali P, Jay B, Paul K, Robert R, Ryan R, Silviu D, Siyu J, Yufan C, Caryn P, David A, Juan Carlos C, Dustin M, Kent P, Zeke W, Victoria O, Iris H, Weibin Z

Online

Tao Y, Adam Z, Ye T, Eric K, Dmitrii N, Don J, Devi A, Jinlong Z, Cameron C,

Cannot make it

KK, Jim N